Circular No. 5610

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 G. Beskin and S. I. Neizvestny, Special Astrophysical Observatory; and

C. Bartolini, A. Guarnieri, and A. Piccioni, University of Bologna, report: “Spectroscopic and photometric observations of GRO J0422+32 at the 6-m telescope show considerable variations of the continuum in the ultraviolet region on a timescale of days. In some spectra, Hβ is seen as a broad absorption with a small emission at its center. Apparently there are several absorption systems in the ranges 380-420 and 460-490 nm. An emission line at 521.8 ± 0.2 nm, possibly due to Fe II, is present. The object becomes strongly ‘violet’ with a practically synchronous brightness increase in U, B, V, and R bands, the U brightness increasing much more than the B. On Sept. 5.918 UT, V = 13.37 ± 0.01, U–B = –0.40 ± 0.02, B–V = +0.15 ± 0.01, V–R = +0.53 ± 0.01. Fast photometry reveals sporadic variability in the B and R bands, in the form of short flashes of timescale 0.2-0.5 s, and relatively smooth variations of duration 10-30 s. Amplitudes of both variations are 10-50 percent.”

FG SAGITTAE R. Gonzalez-Riestra, International Ultraviolet Explorer (IUE) Obser-

vatory, European Space Agency; and H. E. Jorgensen and C. Grunnet, Copenhagen University Observatory, communicate: “Because of the pronounced drop in the visual brightness of FG Sge during the last few weeks (IAUC 5604, 5609), observations of this star were made with the IUE on Sept. 11.6 UT with the long-wavelength spectrograph (195.0-335.0 nm) in the low-dispersion mode (resolution 0.6 nm). The visual magnitude, as measured with the IUE Fine Error Sensor, was ∼ 11.1. The observed flux in the range 270.0-330.0 nm was 2.90 × 10^{−12} erg cm^{−2} s^{−1}, a factor of 3.8 fainter than in the last observations performed in 1991 October, indicating a decrease similar to that seen in the optical. FG Sge is now at the lowest level since it started to be regularly observed with IUE in 1982. The overall spectral shape appears not to be varying. The shape and strength of the Fe II multiplet uv1 near 263.0 nm has varied in a complex way and appears weaker. Infrared observations are needed in order to clarify the origin of the decline in brightness.”

Further visual magnitude estimates: Sept. 10.9 UT, 11.0 (J. Speil, Walbrzych, Poland); 11.98, 11.0 (B. H. Granslo, Fjellhamar, Norway); 13.31, 10.8 (R. Royer, Lakewood, CA); 14.19, 11.4 (D. York, Abiquiu, NM).

1992 September 14 Daniel W. E. Green